OpenStudy (anonymous):

write the sum or difference in the standard form a +bi (squareroot sign 5 -3i) + (-2+ squareroot sign-9)

OpenStudy (shivaniits):

what's ur equation..?? is it\[(\sqrt{5}-3i)+(-2+\sqrt{-9})..??\]

OpenStudy (anonymous):

\[(\sqrt{5} - 3i) + (-2 + \sqrt{-9}) \implies (\sqrt{5} -2) + (-3i + \sqrt{-9})\]

OpenStudy (anonymous):

Remember : \[\sqrt{-9} = i \sqrt{9} = 3i\]

OpenStudy (anonymous):

So: \[(\sqrt{5} -2) + (-3i + 3i) = ??\]

OpenStudy (shivaniits):

i think then \[\sqrt{-9}\]-3i could be written as 0i as \[\sqrt{-9}\]=3i and answer would be only real no \[\sqrt{5}-2\]

OpenStudy (anonymous):

@uri, your girl child is looking very cute in the pic.. ha ha ha..

OpenStudy (uri):

thats a baby boy.

OpenStudy (anonymous):

thanks guys! that helped a lot!

OpenStudy (anonymous):

Oh you have two kids ?? Ha ha ha..

OpenStudy (anonymous):

@sparki1 you are welcome..